# Design principles of the cytotoxic CD8+ T-cell response

### v1

@@ Discussion
We assumed that, although memory formation is essential, its immediate impact on primary effector magnitude is modest: $5-10\%$ of peak effector numbers [38, 1]. Memory “quality” (longevity, proliferative capacity) depends on differentiation timing and the avidity of the selected clones [12, 78, 40, 27]. To isolate how signal integration shapes acute control, we did not explicitly model memory differentiation or recall. An important next step is to extend the framework to the allocation trade-off in which diverting cells to memory lowers primary effector counts but seeds future protection, thereby linking our high-level control design to memory formation and the hosts’ evolutionary survival pressures that could have shaped the effector program.
@@
Key questions follow from our study. Our coarse-grained model elides many details of tissue biology and micro-environment: how do tissue contexts impact the state of resident cells and shift clearance-toxicity trade-offs? Endogenous immune responses are polyclonal [68]: how do the proposed designs shape information flow and selection of dominant clones in polyclonal responses? Do the optimization principles posited here predispose the immune response to certain vulnerabilities, and what mechanisms might have evolved to mask them? Addressing these issues through the lens of cell economics and immunological trade-offs may clarify the roles of chance, contingency, and necessity in the evolution of effector programs.
@@ II Data and code availability
Acknowledgment
This work has been supported by the CAREER award from the National Science Foundation grant 2045054, and the National Institutes of Health MIRA award R35 GM142795. This work is also supported, in part, through the Departments of Physics and Applied Mathematics and the College of Arts and Sciences at the University of Washington. The numerical analyses in this work were completed on Hyak, the UW’s high performance computing cluster, which is funded by the UW student technology fee. This work benefited from discussions during the 2024 program “Interactions and Co-evolution between Viruses and Immune Systems” at the Kavli Institute for theoretical physics (KITP), which is supported by National Science Foundation grant PHY-2309135, and the Gordon and Betty Moore Foundation Grant No. 2919.02.

### v2

@@ Discussion
We assumed that, although memory formation is essential, its immediate impact on primary effector magnitude is modest: $5-10\%$ of peak effector numbers [81, 16]. Memory “quality” (longevity, proliferative capacity) depends on differentiation timing and the avidity of the selected clones [82, 15, 21, 83]. To isolate how signal integration shapes acute control, we did not explicitly model memory differentiation or recall. An important next step is to extend the framework to the allocation trade-off in which diverting cells to memory lowers primary effector counts but seeds future protection, thereby linking our high-level control design to memory formation and the host’s evolutionary survival pressures that could have shaped the effector program.
@@
Key questions follow from our study. Our coarse-grained model elides many details of tissue biology and micro-environment: how do tissue contexts impact the state of resident cells and shift clearance-toxicity trade-offs [88, 89]? Endogenous immune responses are polyclonal [90]: how do the proposed designs shape information flow and selection of dominant clones in polyclonal responses? Do the optimization principles posited here predispose the immune response to certain vulnerabilities, and what mechanisms might have evolved to mask them? Addressing these issues through the lens of cell economics and immunological trade-offs may clarify the roles of chance, contingency, and necessity in the evolution of effector programs.
@@ II Data and code availability
Acknowledgement
This work has been supported by the CAREER award from the National Science Foundation grant 2045054 (A.N., O.A.U., Z.M.), the National Institutes of Health MIRA award R35 GM142795 (A.N.), and the Intramural Research Program of the National Cancer Institute, Project # ZIA BC 012007 (G.A.-B.). This work is also supported, in part, through the Departments of Physics and Applied Mathematics and the College of Arts and Sciences at the University of Washington. The numerical analyses in this work were completed on Hyak, the UW’s high performance computing cluster, which is funded by the UW student technology fee. This work benefited from discussions during the 2023 summer workshop “Statistical Physics and Adaptive Immunity” at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452, and the 2024 program “Interactions and Co-evolution between Viruses and Immune Systems” at the Kavli Institute for theoretical physics (KITP), which is supported by National Science Foundation grant PHY-2309135, and the Gordon and Betty Moore Foundation Grant No. 2919.02.